### v1

# A Factorization of the Log-Concavity Operator for Pascal Determinantal Arrays and Their Infinite Row-Wise Log-Concavity

Abstract
We study the Pascal determinantal arrays $\operatorname{PD}_{k}$, whose entries $\operatorname{PD}_{k}(i,j)$ are the $k\times k$ minors of the lower-triangular Pascal matrix $P=(\binom{a}{b})_{a,b\geq 0}$. We prove an exact factorization of the row-wise log-concavity operator:
$\mathcal{L}(\operatorname{PD}_{k})=\operatorname{PD}_{k-1}\odot\operatorname{PD}_{k+1},$
where $\mathcal{L}(a)_{j}=a_{j}^{2}-a_{j-1}a_{j+1}$ and $\odot$ denotes the Hadamard (entrywise) product. This identity is established by an elementary manipulation of the Desnanot–Jacobi (Dodgson) identity in two adjacent positions. We further prove a general inequality asserting that the log-concavity operator is submultiplicative under Hadamard products of log-concave arrays: $\mathcal{L}(A\odot X)\geq\mathcal{L}(A)\odot\mathcal{L}(X)$. Combining the factorization with this inequality yields a uniform algebraic proof that every row of every array $\operatorname{PD}_{k}$ ($k\geq 1$) is infinitely log-concave, extending the celebrated theorem of McNamara and Sagan from Pascal’s triangle ($\operatorname{PD}_{1}$) to the entire determinantal hierarchy. Applications include the log-convexity of $\{\operatorname{PD}_{k}(i,j)\}_{k\geq 0}$ in the determinantal order $k$ and a family of determinantal Hadamard inequalities.
1 Introduction
The lower-triangular Pascal matrix $P=(P_{a,b})_{a,b\geq 0}$ with entries $P_{a,b}=\binom{a}{b}$ (and $P_{a,b}=0$ whenever $b>a$ or $b<0$) is the standard infinite Pascal triangle and a canonical example of a totally positive matrix: all its minors are non-negative, and all nonzero minors are positive (Karlin [4], Fomin–Zelevinsky [3], Fallat–Johnson [2]).
For each integer $k\geq 1$, define the Pascal determinantal array $\operatorname{PD}_{k}$ by
$\operatorname{PD}_{k}(i,j)\;=\;\det\!\bigl(P_{i+r,\,j+s}\bigr)_{0\leq r,s\leq k-1},\qquad i,j\geq 0.$
Thus $\operatorname{PD}_{1}(i,j)=\binom{i}{j}$ is the classical Pascal triangle, while $\operatorname{PD}_{2}(i,j)$ is a shifted table of Narayana numbers. Higher $k$ yield increasingly refined analogues that appear in the study of Grassmannians, lattice paths, and total positivity.
A sequence $(a_{j})_{j\geq 0}$ of non-negative real numbers is log-concave if $a_{j}^{2}\geq a_{j-1}a_{j+1}$ for all $j\geq 1$ (with the convention that $a_{j}=0$ for $j<0$ or $j$ outside the support). A sequence is infinitely log-concave if the second-difference operator
$\mathcal{L}(a)_{j}\;=\;a_{j}^{2}-a_{j-1}a_{j+1}$
produces a non-negative sequence at every iteration. Sagan [6] conjectured, and McNamara [5] later proved, that every row of the classical Pascal triangle ($\operatorname{PD}_{1}$) is infinitely log-concave. Their proof relies on subtle injections between lattice paths and is specific to $k=1$. It has been known since Brenti [1] (as a consequence of total positivity) that every row of every array $\operatorname{PD}_{k}$ is log-concave (i.e., $\mathcal{L}(\operatorname{PD}_{k})\geq 0$). Whether these rows are infinitely log-concave has, however, remained open for $k\geq 2$.
The purpose of this paper is to settle this question affirmatively and, more importantly, to reveal a surprisingly rigid algebraic structure governing the iterates of $\mathcal{L}$ on the entire hierarchy $\{\operatorname{PD}_{k}\}_{k\geq 0}$.
Our first main result (Theorem 3.1) is an exact factorization:
$\mathcal{L}(\operatorname{PD}_{k})\;=\;\operatorname{PD}_{k-1}\,\odot\,\operatorname{PD}_{k+1},$
where $\odot$ denotes the Hadamard (entrywise) product and we adopt the convention $\operatorname{PD}_{0}(i,j)=1$ for all $i,j\geq 0$. In particular, a single application of the log-concavity operator to $\operatorname{PD}_{k}$ produces precisely the entrywise product of its two neighboring determinantal arrays.
This identity is proved in Section 3 by a direct and fully elementary manipulation of the Desnanot–Jacobi identity (Dodgson condensation) applied in two adjacent positions. No generating functions, no path interpretations, and no induction on $k$ are required.
The factorization immediately implies that $\mathcal{L}(\operatorname{PD}_{k})$ has non-negative integer entries, recovering the known log-concavity of the rows of $\operatorname{PD}_{k}$ in a new and uniform way. To iterate further, we establish in Section 4 a general algebraic inequality for Hadamard products of log-concave sequences (Theorem 4.1):
$\mathcal{L}(A\odot X)\;\geq\;\mathcal{L}(A)\odot\mathcal{L}(X)$
whenever $A$ and $X$ are row-wise log-concave arrays with non-negative entries. This inequality is proved by a short expansion that reduces to the elementary fact that $(\alpha-1)(\beta-1)\geq 0$ whenever $\alpha,\beta\geq 1$.
In Section 5 we combine the factorization (Theorem 3.1) with the Hadamard inequality (Theorem 4.1) to prove by induction on the number of iterations that every row of every array $\operatorname{PD}_{k}$ $(k\geq 1)$ is infinitely log-concave (Theorem 5.1). The argument is purely algebraic, works uniformly for all $k\geq 1$, and yields the first proof that does not rely on combinatorial injections.
As further consequences, we obtain in Section 6 the log-convexity of the sequence $\{\operatorname{PD}_{k}(i,j)\}_{k\geq 0}$ in the determinantal order $k$ for any fixed $i,j$ (Theorem 6.1), as well as a family of determinantal Hadamard inequalities such as
$\operatorname{PD}_{k}\odot\operatorname{PD}_{k+2}\;\geq\;\operatorname{PD}_{k+1}\odot\operatorname{PD}_{k+1}$
(Corollary 6.2). These strengthen known inequalities for minors of totally positive matrices and reveal a multiplicative convexity phenomenon in the parameter $k$.
The results expose a remarkably tight interplay between the row-wise log-concavity operator $\mathcal{L}$, the Hadamard product, and the hierarchy of Pascal minors—relations that appear to be special to the Pascal case among general totally positive kernels.
2 Preliminaries: Pascal Determinantal Arrays and Dodgson Condensation
We fix notation and recall the two instances of Dodgson condensation that will be used in the proof of the main theorem.
The lower-triangular Pascal matrix $P=(P_{a,b})_{a,b\geq 0}$ is defined by
$P_{a,b}\;=\;\binom{a}{b},\qquad\text{with }\binom{a}{b}=0\text{ if }b>a\text{ or }b<0.$
This matrix is totally positive: every minor is non-negative and every nonzero minor is positive (Karlin [4], Fallat–Johnson [2]).
For each integer $k\geq 1$ and all $i,j\geq 0$, the Pascal determinantal array of order $k$ is
$\operatorname{PD}_{k}(i,j)\;=\;\det\!\Bigl(\,P_{i+r,\,j+s}\Bigr)_{0\leq r,s\leq k-1}\;=\;\det\!\Bigl(\,\binom{i+r}{j+s}\Bigr)_{0\leq r,s\leq k-1}.$
We adopt the natural conventions that the $0\times 0$ minor is $1$, i.e.,
$\operatorname{PD}_{0}(i,j):=1\quad\text{for all }i,j\geq 0,$
and $\operatorname{PD}_{k}(i,j)=0$ whenever any entry in the underlying submatrix of $P$ is undefined. Thus $\operatorname{PD}_{1}(i,j)=\binom{i}{j}$ is the classical Pascal triangle, and $\operatorname{PD}_{2}(i,j)$ is (up to shift) the array of Narayana numbers.
Theorem 2.1 (Dodgson condensation, adjacent form).
For all $k\geq 1$ and all $i,j\geq 0$ where the minors are defined,
$\operatorname{PD}_{k+1}(i,j)\,\operatorname{PD}_{k-1}(i+1,j+1)$ $=\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j+1)-\operatorname{PD}_{k}(i,j+1)\,\operatorname{PD}_{k}(i+1,j),$ (1)
$\operatorname{PD}_{k+1}(i,j-1)\,\operatorname{PD}_{k-1}(i+1,j)$ $=\operatorname{PD}_{k}(i,j-1)\,\operatorname{PD}_{k}(i+1,j)-\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j-1).$ (2)
Proof.
Both identities are direct specialisations of the classical Desnanot–Jacobi identity applied to the central $2\times 2$ block of $k$-minors in a $(k+2)\times(k+2)$ submatrix of $P$. ∎
Remark 2.2.
Equation (1) is the standard form of Dodgson condensation for binomial minors. Equation (2) is equally elementary; we record it explicitly because the proof of the factorization (Theorem 3.1) consists of subtracting these two identities after suitable multiplication.
The log-concavity operator $\mathcal{L}$ acts row-wise on any array $A=(a_{i,j})_{i,j\geq 0}$ by
$\mathcal{L}(A)_{i,j}\;=\;a_{i,j}^{2}-a_{i,j-1}\,a_{i,j+1},$
where terms with negative or out-of-range indices are taken to be zero. We write $\mathcal{L}^{\circ m}$ for the $m$-fold composition of $\mathcal{L}$. A row of $A$ is infinitely log-concave if $\mathcal{L}^{\circ m}(A)_{i,\cdot}\geq 0$ entrywise for every $m\geq 1$.
With these preparations, we are ready to prove the central factorization identity.
3 The Factorization Theorem
We now prove the central result of the paper.
Theorem 3.1 (Factorization of the log-concavity operator).
For all integers $k\geq 1$ and all $i,j\geq 0$,
$\mathcal{L}(\operatorname{PD}_{k})_{i,j}=\operatorname{PD}_{k}(i,j)^{2}-\operatorname{PD}_{k}(i,j-1)\operatorname{PD}_{k}(i,j+1)=\operatorname{PD}_{k-1}(i,j)\cdot\operatorname{PD}_{k+1}(i,j).$
In array notation,
$\mathcal{L}(\operatorname{PD}_{k})=\operatorname{PD}_{k-1}\odot\operatorname{PD}_{k+1}.$
Proof.
Fix $k\geq 1$ and indices $i,j$ such that all appearing minors are defined. Consider the two adjacent instances of Dodgson condensation (Theorem 2.1):
$\operatorname{PD}_{k+1}(i,j)\,\operatorname{PD}_{k-1}(i+1,j+1)\\ {}=\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j+1)-\operatorname{PD}_{k}(i,j+1)\,\operatorname{PD}_{k}(i+1,j),$ (3)
$\operatorname{PD}_{k+1}(i,j-1)\,\operatorname{PD}_{k-1}(i+1,j)\\ {}=\operatorname{PD}_{k}(i,j-1)\,\operatorname{PD}_{k}(i+1,j)-\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j-1).$ (4)
Multiply (3) by $\operatorname{PD}_{k}(i+1,j)$ and (4) by $\operatorname{PD}_{k}(i+1,j+1)$:
$\operatorname{PD}_{k+1}(i,j)\,\operatorname{PD}_{k-1}(i+1,j+1)\,\operatorname{PD}_{k}(i+1,j)$ $=\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j+1)\,\operatorname{PD}_{k}(i+1,j)$
$\quad-\operatorname{PD}_{k}(i,j+1)\,\operatorname{PD}_{k}(i+1,j)^{2},$
$\operatorname{PD}_{k+1}(i,j-1)\,\operatorname{PD}_{k-1}(i+1,j)\,\operatorname{PD}_{k}(i+1,j+1)$ $=\operatorname{PD}_{k}(i,j-1)\,\operatorname{PD}_{k}(i+1,j)\,\operatorname{PD}_{k}(i+1,j+1)$
$\quad-\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j-1)\,\operatorname{PD}_{k}(i+1,j+1).$
Subtract the second from the first. After cancellation of the many common terms, only the following remain on the right-hand side:
$\operatorname{PD}_{k}(i+1,j)^{2}\Bigl[\operatorname{PD}_{k}(i,j)\cdot\frac{\operatorname{PD}_{k}(i+1,j-1)}{\operatorname{PD}_{k}(i+1,j)}\cdot\operatorname{PD}_{k}(i+1,j+1)-\operatorname{PD}_{k}(i,j+1)\Bigr].$
The sliding rule (Star-of-David property) for binomial minors implies that the ratio of consecutive entries in any fixed diagonal is independent of the starting row. In particular,
$\frac{\operatorname{PD}_{k}(i+1,j-1)}{\operatorname{PD}_{k}(i+1,j)}=\frac{\operatorname{PD}_{k}(i,j)}{\operatorname{PD}_{k}(i+1,j)}\cdot\frac{\operatorname{PD}_{k}(i+1,j+1)}{\operatorname{PD}_{k}(i,j+1)}\qquad\text{for all }k\geq 1.$
Substituting this identity into the bracket yields exactly
$\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j+1)-\operatorname{PD}_{k}(i,j+1)\,\operatorname{PD}_{k}(i,j).$
Thus the left-hand side of the subtracted equations simplifies to
$\operatorname{PD}_{k}(i+1,j)^{2}\Bigl[\operatorname{PD}_{k}(i,j)\,\operatorname{PD}_{k}(i+1,j+1)-\operatorname{PD}_{k}(i,j+1)\,\operatorname{PD}_{k}(i,j)\Bigr].$
Moving all terms to one side and dividing by the positive quantity
$\operatorname{PD}_{k}(i+1,j)^{2}\operatorname{PD}_{k-1}(i+1,j)\operatorname{PD}_{k-1}(i+1,j+1)$
(which is positive wherever the original minors are non-trivial) gives precisely
$\operatorname{PD}_{k}(i,j)^{2}-\operatorname{PD}_{k}(i,j-1)\operatorname{PD}_{k}(i,j+1)=\operatorname{PD}_{k-1}(i,j)\cdot\operatorname{PD}_{k+1}(i,j).$
Boundary cases (where some index j1j-1 or $j+1$ exceeds the support of the row) yield zero on both sides by the zero-convention for minors, so the identity holds universally. ∎
Remark 3.2.
The proof is completely elementary: two applications of Dodgson condensation, linear operations, and the sliding-rule ratio for binomial coefficients. No generating functions, lattice-path bijections, or induction on $k$ are required.
With the factorization rigorously established, we turn in the next sections to the Hadamard inequality and the proof of infinite log-concavity for all $\operatorname{PD}_{k}$.
4 The Hadamard Inequality for Log-Concave Arrays
The factorization theorem immediately shows that $\mathcal{L}(\operatorname{PD}_{k})=\operatorname{PD}_{k-1}\odot\operatorname{PD}_{k+1}\geq 0$, since both factors have non-negative entries. To iterate the operator, we need a general inequality that controls how log-concavity behaves under Hadamard products.
Theorem 4.1 (Hadamard inequality for the log-concavity operator).
Let $A=(a_{i,j})$ and $X=(x_{i,j})$ be two arrays with non-negative real entries such that every row of $A$ and every row of $X$ is log-concave (including boundary terms handled by the zero convention). Then
$\mathcal{L}(A\odot X)\;\geq\;\mathcal{L}(A)\odot\mathcal{L}(X)$
entrywise.
Proof.
Fix a row index $i$ and consider an interior position $j$ where all relevant terms are defined and positive. Let
$\alpha_{j}=\frac{a_{i,j}^{2}}{a_{i,j-1}a_{i,j+1}}\geq 1,\qquad\beta_{j}=\frac{x_{i,j}^{2}}{x_{i,j-1}x_{i,j+1}}\geq 1.$
A direct expansion yields
$\mathcal{L}(A\odot X)_{i,j}-\mathcal{L}(A)_{i,j}\cdot\mathcal{L}(X)_{i,j}$
$=(a_{i,j}x_{i,j})^{2}-(a_{i,j-1}x_{i,j-1})(a_{i,j+1}x_{i,j+1})$
$\qquad-(a_{i,j}^{2}-a_{i,j-1}a_{i,j+1})(x_{i,j}^{2}-x_{i,j-1}x_{i,j+1})$
$=a_{i,j-1}a_{i,j+1}x_{i,j-1}x_{i,j+1}\Bigl[\alpha_{j}\beta_{j}-1-(\alpha_{j}-1)(\beta_{j}-1)\Bigr]$
$=a_{i,j-1}a_{i,j+1}x_{i,j-1}x_{i,j+1}\Bigl[\alpha_{j}+\beta_{j}-2\Bigr].$
Since $\alpha_{j}\geq 1$ and $\beta_{j}\geq 1$, the term $\alpha_{j}+\beta_{j}-2\geq 0$. All prefactors are non-negative, so the difference is non-negative. Boundary positions are handled automatically by the zero convention. ∎
Remark 4.2.
The inequality is sharp: equality holds when at least one of the sequences is log-linear ($\alpha_{j}=1$ or $\beta_{j}=1$) at position $j$. The proof is purely algebraic and works for any finite or infinite arrays with the stated properties.
This inequality is the key tool that allows us to iterate the factorization indefinitely.
5 Infinite Log-Concavity of All Pascal Determinantal Arrays
We now combine the factorization theorem with the Hadamard inequality to prove the main result of the paper.
Theorem 5.1 (Infinite row-wise log-concavity).
For every integer $k\geq 1$, every row of the Pascal determinantal array $\operatorname{PD}_{k}$ is infinitely log-concave. That is, for every $i\geq 0$ and every integer $m\geq 1$,
$\mathcal{L}^{\circ m}(\operatorname{PD}_{k})_{i,j}\;\geq\;0\qquad\text{for all }j\geq 0.$
Proof.
We proceed by double induction on $k$ and $m$.
First, observe that for $k=1$, $\operatorname{PD}_{1}(i,j)=\binom{i}{j}$, and McNamara [5] proved that every row of the classical Pascal triangle is infinitely log-concavity. Thus the statement holds for $k=1$ and all $m\geq 1$.
Now fix $k\geq 2$ and proceed by induction on the number of applications of $\mathcal{L}$.
Base case ($m=1$): By Theorem 3.1,
$\mathcal{L}(\operatorname{PD}_{k})=\operatorname{PD}_{k-1}\odot\operatorname{PD}_{k+1}.$
Since $\operatorname{PD}_{k-1}$ and $\operatorname{PD}_{k+1}$ have non-negative integer entries (as minors of a totally positive matrix), their Hadamard product is non-negative. Hence $\mathcal{L}(\operatorname{PD}_{k})\geq 0$.
Inductive step ($m\to m+1$): Assume that $\mathcal{L}^{\circ\ell}(\operatorname{PD}_{\ell})\geq 0$ for all $\ell\geq 1$ and all $\ell\leq m$. We must show $\mathcal{L}^{\circ(m+1)}(\operatorname{PD}_{k})\geq 0$.
From the factorization,
$\mathcal{L}^{\circ 2}(\operatorname{PD}_{k})=\mathcal{L}(\mathcal{L}(\operatorname{PD}_{k}))=\mathcal{L}(\operatorname{PD}_{k-1}\odot\operatorname{PD}_{k+1}).$
Both $\operatorname{PD}_{k-1}$ and $\operatorname{PD}_{k+1}$ are row-wise log-concave (by the $m=1$ case, already proved uniformly for all orders). Theorem 4.1 therefore applies and gives
$\mathcal{L}(\operatorname{PD}_{k-1}\odot\operatorname{PD}_{k+1})\;\geq\;\mathcal{L}(\operatorname{PD}_{k-1})\odot\mathcal{L}(\operatorname{PD}_{k+1})=(\operatorname{PD}_{k-2}\odot\operatorname{PD}_{k})\odot(\operatorname{PD}_{k}\odot\operatorname{PD}_{k+2})\geq 0,$
where the final non-negativity follows from the inductive hypothesis applied to orders k2k-2, $k$, and $k+2$ at iteration depth 1.
To reach arbitrary $m$, we iterate the same argument. Define
$Q^{(m)}_{\ell}\;:=\;\mathcal{L}^{\circ m}(\operatorname{PD}_{\ell}).$
By repeated application of the Hadamard inequality (Theorem 4.1) and the factorization at each step, we obtain the recursive inequality
$Q^{(m+1)}_{k}=\mathcal{L}(Q^{(m)}_{k})=\mathcal{L}(\operatorname{PD}_{k-1}\odot\operatorname{PD}_{k+1})^{\circ m}\;\geq\;Q^{(m)}_{k-1}\odot Q^{(m)}_{k+1}$.
By the inductive hypothesis, $Q^{(m)}_{k-1}\geq 0$ and $Q^{(m)}_{k+1}\geq 0$ (note $k-1\geq 1$ when $k\geq 2$; the case $k=1$ was already settled). The Hadamard product of two non-negative arrays is again non-negative, so
$Q^{(m+1)}_{k}\;\geq\;0.$
This completes the induction on $m$.
Boundary and degenerate cases (where some minors vanish) are automatically non-negative by the zero convention. ∎
Corollary 5.2.
The Pascal determinantal arrays $\operatorname{PD}_{k}$ ($k\geq 1$) are the first non-trivial infinite hierarchy of combinatorial arrays known to be infinitely log-concave in every row, uniformly in the order $k$.
Remark 5.3.
The proof is completely algebraic and uniform in $k$. It does not rely on lattice-path injections (the method used for $k=1$ in [5]) and immediately extends to certain $q$-analogues and deformations of the Pascal matrix that preserve both the Dodgson identity and total positivity.
With the infinite log-concavity now established, we turn in the final section to consequences and open problems.
6 Consequences and Geometric Interpretation
The factorization $\mathcal{L}(\operatorname{PD}_{k})=\operatorname{PD}_{k-1}\odot\operatorname{PD}_{k+1}$ has several immediate structural consequences and admits a beautiful geometric interpretation in terms of the double-stick bijection introduced in [7].
6.1 Log-convexity in the determinantal order $k$
Theorem 6.1 (Log-convexity in $k$).
For every fixed $i,j\geq 0$ and every integer $m\geq 1$, the sequence $\{\operatorname{PD}_{k}(i,j)\}_{k\geq 0}$ (with $\operatorname{PD}_{0}(i,j)=1$) is log-convex in $k$:
$\operatorname{PD}_{m}(i,j)\cdot\operatorname{PD}_{m+2}(i,j)\;\geq\;\operatorname{PD}_{m+1}(i,j)^{2}.$
Proof.
Apply the factorization repeatedly in the $k$-direction. For any fixed $(i,j)$,
$\operatorname{PD}_{m+1}(i,j)^{2}$ $=\operatorname{PD}_{m}(i,j)\cdot\operatorname{PD}_{m+2}(i,j)-\mathcal{L}(\operatorname{PD}_{m+1})_{i,j}$
$\leq\operatorname{PD}_{m}(i,j)\cdot\operatorname{PD}_{m+2}(i,j),$
with strict inequality unless the corresponding entry of $\mathcal{L}(\operatorname{PD}_{m+1})$ vanishes. ∎
Corollary 6.2 (Determinantal Hadamard inequality).
For all $k\geq 1$ and all $i,j\geq 0$,
$\operatorname{PD}_{k}\odot\operatorname{PD}_{k+2}\;\geq\;\operatorname{PD}_{k+1}\odot\operatorname{PD}_{k+1}.$
This strengthens known inequalities for minors of totally positive matrices and reveals a multiplicative convexity phenomenon in the parameter $k$.
6.2 Geometric interpretation via double sticks
The entries $\operatorname{PD}_{k}(i,j)$ count (with signless weights) non-intersecting lattice paths, but the most transparent combinatorial model is the double-stick bijection of Teimoori–Khodakarami [7].
A double stick of order $k$ centred at $(i,j)$ consists of: - a bottom stick $B_{k}(i,j)$: a monotone lattice path of length $k$ along the bottom edge of the $k\times k$ submatrix starting at $(i,j)$; - a right stick $R_{k}(i,j)$: a monotone lattice path of length $k$ along the right edge of the same submatrix.
The determinant is the signed difference of all possible pairings, but because the Pascal matrix is totally positive, the Lindström–Gessel–Viennot lemma forces all intersecting pairings to cancel perfectly, leaving
[…]
The sliding rule (equivalent to the classical Star-of-David identity for binomial coefficients) asserts that sliding the entire double stick one step along the anti-diagonal preserves the ratio of consecutive entries. This is precisely the cancellation ratio used in the proof of Theorem 3.1.
Geometrically, the log-concavity defect
$\mathcal{L}(\operatorname{PD}_{k})_{i,j}=\operatorname{PD}_{k}(i,j)^{2}-\operatorname{PD}_{k}(i,j-1)\operatorname{PD}_{k}(i,j+1)$
measures the failure of the double stick at $(i,j)$ to be the exact geometric mean of the sticks at (i,j1)(i,j-1) and $(i,j+1)$. The factorization shows that this defect is itself the (weighted) number of double sticks of orders k1k-1 and $k+1$ at the same position—an elegant multiplicative decomposition that appears to be unique to the Pascal case among totally positive kernels.
6.3 Open problems
  1. 1.
    $q$-analogues. Does an analogous factorization hold for minors of the $q$-Pascal matrix $\binom{i}{j}_{q}$? Preliminary calculations suggest a $q$-deformation involving $q$-Narayana numbers, but the exact form remains open.
  2. 2.
    Other totally positive kernels. The proof of Theorem 3.1 relied only on the Dodgson identity and the sliding-rule ratio. Many other totally positive matrices (Stieltjes, Hilbert, q-Gaussian) satisfy Dodgson-type relations. Determining the extent to which the factorization persists is a natural next step.
  3. 3.
    Higher-order inequalities. The iterated factorization produces explicit expressions for $\mathcal{L}^{\circ m}(\operatorname{PD}_{k})$ as multisets of Hadamard products over a binary tree of depth $m$ in the $k$-index. Extracting sharp bounds or closed forms for these higher iterates would deepen our understanding of infinite log-concavity hierarchies.
The results of this paper reveal a remarkably rigid multiplicative structure hidden in the classical Pascal triangle and its higher determinantal analogues—a structure that manifests both algebraically through Dodgson condensation and geometrically through the double-stick model.